OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = -8,748.

OpenStudy (anonymous):

Any idea?? Hint \(a_n = a_1*r^{n-1}\) where r is ratio

OpenStudy (anonymous):

no not really sure @OOOPS

OpenStudy (anonymous):

a2 = a1*r a3 = a2*r= a1*r^2 a4 = a3*r = a1*r^3 and so on, can you give me the expression of a8=?

OpenStudy (anonymous):

what is the ratio though?

OpenStudy (anonymous):

just let r for it, we have to find it out from a8

OpenStudy (anonymous):

dosnt in =8748

OpenStudy (anonymous):

I want you to represent a8 = a1* r ^??